Marst Corporation's budgeted production in units and budgeted raw materials purchases over the next three months are given below:

January February March
Budgeted production (in units) 78,000 ? 80,000
Budgeted raw materials purchases (in pounds) 174,600 200,600 256,600


Two pounds of raw materials are required to produce one unit of product. The company wants raw materials on hand at the end of each month equal to 30% of the following month's production needs. The company is expected to have 28,000 pounds of raw materials on hand on January 1. Budgeted production for February should be:

rev: 10_27_2016_QC_CS-67319

48,000 units

109,000 units

96,000 units

152,600 units

question 2)

Paradise Corporation budgets on an annual basis for its fiscal year. The following beginning and ending inventory levels (in units) are planned for next year.

Beginning Inventory Ending Inventory
Raw material* 63,800 60,600
Finished goods 84,000 53,000


* Three pounds of raw material are needed to produce each unit of finished product.

If Paradise Corporation plans to sell 495,000 units during next year, the number of units it would have to manufacture during the year would be:

Noreen 4e Recheck 2017-22-03

494,000 units

454,000 units

464,000 units

524,000 units
Answer #1
1
Let Budgeted production for February be X units
200600 = 2X +(80000*2*30%)-(2X*30%)
200600 = 2X + 48000 - 0.6X
200600-48000 = 1.4X
152600 = 1.4X
X = 152600/1.4= 109,000 units
Option B 109,000 units is correct
2
Budgeted production = Budgeted sales+Finished goods Ending Inventory-Finished goods Beginning Inventory
Budgeted production = 495000+53000-84000= 464,000 units
Option C 464,000 units is correct